Please help me anyone please thank you.
Problem is 6.161616 and repeate 1616, 3.4125125 repeate 125 I need find ratio also here is formula a1 all over (1-ratio)

Answer by Alan3354(69443)

Problem is 6.161616 and repeate 1616,
n = 0.161616...
100n = 16.161616...
+1n = 0.161616...
-------------------- Subtract
99n = 16
n = 16/99
--> 6.161616,,, = 6 + 16/99 or 610/99
-----------------
3.4125125 repeate 125 I need find ratio
n = 0.0125125...
1000n = 12.5125125...
Sub n = 12.5
999n = 12.5n
n = 12.5/99 = 25/198
--> 3.4 + 25/1998 = 34/10 + 25/1998
= 34091/9990